OpenStudy (anonymous):

Find the minimum, maximum, range, mean, and median for the data set. 2.4, 2.6, 7.2, 5.5, 4.1, 5.5, 6.1, 6.1 A. minimum: 2.4 maximum: 7.2 range: 4.8 mean: 5.5 median: 4.9 B. minimum: 2.4 maximum: 7.2 range: 9.6 mean: 4.9 median: 5.5 C. minimum: 7.2 maximum: 2.4 range: 4.8 mean: 4.9 median: 5.5 D. minimum: 2.4 maximum: 7.2 range: 4.8 mean: 4.9 median: 5.5

OpenStudy (anonymous):

@Sam_Aka_Sara @unicorn_princess @SithsAndGiggles @lolz? @amistre64 @Quxxn @Luigi0210 @justuu @KenLJW @whpalmer4 @grex10 @BloomLocke367 @Xaze @Awolflover1 @math&ing001 @darkb0nebeauty @Concentrationalizing @zoefashionista @gabby_2624 @adajiamcneal @EWack @ellenaschaefer @foresee123 @FloppyRawrRawr @rockstar0765 @youngman @yowasup @yolomcswagginsggg @briggs

OpenStudy (luigi0210):

The mass tagging is real

OpenStudy (anonymous):

@Luigi0210 I dont even know you guys i was just doing the @ and clicking

OpenStudy (anonymous):

lol im not good at this but medal me lool :D

OpenStudy (anonymous):

Nty

OpenStudy (anonymous):

(-_-)

OpenStudy (anonymous):

I HAVE NONE LOL ILY!

OpenStudy (anonymous):

hehe :)

OpenStudy (yolomcswagginsggg):

first arrange the data set from higher value to the lower value... the lowest value is the minimum highest value is the maximum and the middle term that u get when u arrange them in the above order is the median the difference between the highest value and the lowest is the range.. add them altogether and divide them by the no.of terms..thats the mean can u do it ?

OpenStudy (yolomcswagginsggg):

2.4 2.6 4.1 5.5 5.5 6.1 6.1 7.2

OpenStudy (yolomcswagginsggg):

7.2 is the max.

OpenStudy (yolomcswagginsggg):

2.4 is the least

OpenStudy (yolomcswagginsggg):

5.5, and 5.5 are in the middle, their the median.

OpenStudy (yolomcswagginsggg):

median = (5.5 plus 5.5)/2 = 5.5

OpenStudy (yolomcswagginsggg):

range = max -min = 7.2 - 2.4 = 4.8

OpenStudy (yolomcswagginsggg):

2.4 2.6 4.1 5.5 5.5 6.1 6.1 7.2 mean = 2.4 + 2.6+4.1+5.5+5.5+6.1+6.1+7.2 8 (since there r 8 terms) =4.93 which is 4.9

OpenStudy (yolomcswagginsggg):

In other words, its D!